# Ground Perturbation Detection via Lower-Limb Kinematic States During Locomotion

### v2

@@ IV Discussion
his paper, we proposed a novel perturbation detection model, which utilizes 16 lower-limb kinematic states and their deviations from steady-state gait dynamics. We established a local coordinate system to track relative positions to avoid misinterpreting natural drift in the state during walking. Using a data-driven approach, we further optimized a single threshold value to detect ground perturbations, $\varphi$. We hypothesized that our kinematic state-based detection model would outperform the WBAM-based model under varying perturbation conditions across subjects. To test this hypothesis, we evaluated the performance of our model and the WBAM-based benchmark in detecting trip- and slip-type perturbations. Our study results corroborate our initial hypothesis that the ground perturbation detection model using kinematic states detected perturbations more accurately than the WBAM-based model. In the open-source dataset analysis, our model showed a 41% improvement in detection accuracy compared to that of the WBAM-based model (Fig. 6). However, contrary to our initial hypothesis, the WBAM-based model outperformed our model in the detection delay, demonstrating a faster response on average by 11% earlier within the gait cycle. However, it is important to note that this delay was only calculated in instances of successful perturbation detections and may not be representative of the remaining 59% of perturbations that the WBAM-based model failed to detect. Furthermore, since this was an offline analysis, it did not consider the time required to calculate WBAM.
@@
In our extended perturbation pilot experiment, our model demonstrated an improvement in detection accuracy of 10% compared to its performance on the open-source dataset. Furthermore, our model detected comparatively 5% faster within the gait cycle. These results support our sub-hypothesis that increasing the number of steady-state gait data (for nominal state trajectories) would enhance the performance of our model. The generalizability of our model across datasets showcased the potential benefit of integrating our approach into lower-limb exoskeleton control. The average stride duration across subjects in our pilot experiment was 1006.9$\pm$8.3 ms. Biomechanics literature states that older adults can require up to 469 ms to complete a stabilizing step in response to perturbations. Given that the detection timing of our model was 231 ms, our result indicated that the detection was early enough for an exoskeleton to respond (with assistance) within a single stride [4, 33, 3]. Moreover, the generalizability of our model and its consistent performance across subjects showed that the system can potentially be deployed to exoskeletons with minimal user-specific tuning [34]. This subject-independent characteristic of our model was highlighted with our consistent result across two distinct datasets.
@@
While our model was capable of operating in real-time, gap fill and minimizing computation (though its minimal compared to the WBAM approach), when using live-stream marker data, needs further refinement for a real-time deployment. In future work, we aim to make our model fully autonomous by tracking kinematic states using on-board wearable sensors (e.g,. inertial measurement units) from an exoskeleton. Furthermore, we plan to reduce the amount of gait cycle data required to determine the initial nominal state trajectories (current system requiring 10 gait cycles), allowing for more dynamic locomotion [21].
@@
Lastly, we plan to extend our model to determine both the direction and magnitude of each perturbation. The instantaneous sign and magnitude of each $\alpha$ value can provide an estimate of the direction in which the individual’s feet and COM were perturbed as well as the relative magnitude of the perturbation (Fig. 3D). We will evaluate the accuracy of these estimations and explore their potential to enhance the precision of perturbation responses in lower-limb exoskeletons. Moreover, we will evaluate our model on a larger cohort and under diverse ground perturbation conditions to improve the generalizability of our model.
@@ V Conclusion
We developed a novel perturbation detection model by tracking lower-limb kinematic states and their deviations from steady-state gait dynamics during walking. Our model employed a weighted sum of state deviations and used a data-driven threshold to detect ground perturbations. In a pilot human subject experiment, our model achieved 97% detection accuracy with a delay of 23% within a gait cycle after a perturbation onset, outperforming the benchmark WBAM-based model. These results highlight the potential of our model to enable lower-limb exoskeletons to provide fast and precise balance support, particularly for older adult users.

### v3

@@ IV Discussion
In this paper, we proposed a novel perturbation detection model, which utilizes 16 lower-limb kinematic states and their deviations from steady-state gait dynamics. We established a local coordinate system to track relative positions to avoid misinterpreting natural drift in the state during walking. Using a data-driven approach, we further optimized a single threshold value to detect ground perturbations, $\varphi$. We hypothesized that our kinematic state-based detection model would outperform the WBAM-based model under varying perturbation conditions across subjects. To test this hypothesis, we evaluated the performance of our model and the WBAM-based benchmark in detecting trip- and slip-type perturbations. Our study results corroborate our initial hypothesis that the ground perturbation detection model using kinematic states detected perturbations more accurately than the WBAM-based model. In the open-source dataset analysis, our model showed a 41% improvement in detection accuracy compared to that of the WBAM-based model (Fig. 6). However, contrary to our initial hypothesis, the WBAM-based model outperformed our model in the detection delay, demonstrating a faster response on average by 11% earlier within the gait cycle. However, it is important to note that this delay was only calculated in instances of successful perturbation detections and may not be representative of the remaining 54% of perturbations that the WBAM-based model failed to detect. Furthermore, since this was an offline analysis, it did not consider the time required to calculate WBAM.
@@
In our extended perturbation pilot experiment, our model demonstrated an improvement in detection accuracy compared to its performance on the open-source dataset, achieving 47.7% higher accuracy than the WBAM-based model. Furthermore, our model detected 5% faster within the gait cycle compared to its performance on the open-source dataset. The generalizability of our model across datasets showcased the potential benefit of integrating our approach into lower-limb exoskeleton control. The average stride duration across subjects in our pilot experiment was 1006.9$\pm$8.3 ms. Biomechanics literature states that older adults can require up to 469 ms to complete a stabilizing step in response to perturbations. Given that the detection timing of our model was 231 ms, our result indicated that the detection was early enough for an exoskeleton to respond (with assistance) within a single stride [4, 32, 3]. Moreover, the generalizability of our model and its consistent performance across subjects showed that the system can potentially be deployed to exoskeletons with minimal user-specific tuning [33]. This subject-independent characteristic of our model was highlighted with our consistent result across two distinct datasets.
@@
While our model was capable of operating in real-time, gap fill and minimizing computation (though minimal compared to the WBAM approach) when using live-stream marker data needs further refinement for a real-time deployment. In future work, we aim to make our model fully autonomous by tracking kinematic states using on-board wearable sensors (e.g,. inertial measurement units) from an exoskeleton. Furthermore, we plan to reduce the amount of gait cycle data required to determine the initial nominal state trajectories (current system requiring 10 gait cycles), allowing for more dynamic locomotion [26]. Current gait literature states that 2-3 strides are sufficient to capture the natural variance of healthy gait [34], however, further analysis is required to determine the minimum number of strides for our model application, specifically within older adult populations.
@@
Lastly, we plan to extend our model to determine both the direction and magnitude of each perturbation. The instantaneous sign and magnitude of each $\alpha$ value can provide an estimate of the direction in which the individual’s feet and COM were perturbed as well as the relative magnitude of the perturbation (Fig. 3D). We will evaluate the accuracy of these estimations and explore their potential to enhance the precision of perturbation responses in lower-limb exoskeletons. Moreover, we will evaluate our model under diverse ground perturbation conditions and subject demographics to improve the generalizability of our model. Although our model was trained on healthy young subjects, its performance at the same walking speed is likely comparable in healthy older adults. Given their longer reaction times, our method could be especially valuable for assessing reactive balance in older adults [35]. However, additional data collection from older adult subjects is needed to evaluate the model’s performance in this target population.
@@ V Conclusion
We developed a novel perturbation detection model by tracking lower-limb kinematic states and their deviations from steady-state gait dynamics during walking. Our model employed a weighted sum of state deviations and used a data-driven threshold to detect ground perturbations. In a pilot human subject experiment, our model achieved 98% detection accuracy with a delay of 23% within a gait cycle after a perturbation onset, outperforming the benchmark WBAM-based model. These results highlight the potential of our model to enable lower-limb exoskeletons to provide fast and precise balance support, particularly for older adult users.